A clinic is considering the possibility of two new purchases: new MRI equipment and new biopsy equipment. Each project requires an investment of $434,200. The expected life for each is five years with no expected salvage value. The net cash inflows associated with the two independent projects are as follows:


Year

MRI Equipment

Biopsy Equipment
1 $183,000         $60,000        
2 97,000         48,000        
3 166,000         99,000        
4 98,000         182,000        
5 49,000         242,000        

The present value tables provided in Exhibit 19B.1 and Exhibit 19B.2 must be used to solve the following problems.

Required:

Compute the net present value of each project, assuming a required rate of 12 percent. If the NPV is negative, enter your answer as a negative value.

NPV
MRI equipment $
Biopsy equipment $
Answer #1

Answer-Net present value-MRI Equipment =$14770.

Explanation-

Calculation of Investment's Net Present Value
MRI Equipment
Net Cash Flows $ (a) Present Value of 1 at 12% (b) Present Value of cash flows (c=a*b) $
Year 1 183000 0.8929 163401
Year 2 97000 0.7972 77328
Year 3 166000 0.7118 118159
Year 4 98000 0.6355 62279
Year 5 49000 0.5674 27803
Totals
Total present value of cash inflow (a) 448970
Total cash outflow (b) -434200 1 -434200
Net Present Value $ (c=a-b) 14770

Net present value-Blopsy Equipment =($18920).

Explanation-

Calculation of Investment's Net Present Value
Blopsy Equipment
Net Cash Flows $ (a) Present Value of 1 at 12% (b) Present Value of cash flows (c=a*b) $
Year 1 60000 0.8929 53574
Year 2 48000 0.7972 38266
Year 3 99000 0.7118 70468
Year 4 182000 0.6355 115661
Year 5 242000 0.5674 137311
Totals
Total present value of cash inflow (a) 415280
Total cash outflow (b) -434200 1 -434200
Net Present Value $ (c=a-b) -18920